A constant current was passed through a solution of AuCl4 ions between gold electrolyte .After a period of 10 minutes,the cathode increased in weight by 1.314g.how much charge was passed and what is the current?[Au=196.97]
1
Expert's answer
2019-07-31T01:48:59-0400

We take advantage of the united Faraday law:

"m = {1 \over F} {M \over z} It"

(where "F = 96485; z = 4; t = 10 min = 600sec")


"I = {Fmz \over Mt}"

"I = {96485*1.314*4 \over 600*196.97} = 4.291A"

Find current:

"q = It""q = 4.291*600 = 2574.631C"

Answer: I = 4.291 A; q = 2574.631 C.